1. What is the percent composition by mass Carbon in carbon monoxide (CO)?


2. Calculate the percent composition by mass of Oxygen in

baking soda (NaHCO​3)​ .


3. Calculate the percent composition by mass of carbon in sucrose (C​12 H​22 O​11)​ ?


4. In which compound is the percent by mass of sodium the greatest?

a. Na​2S​O4​ b. Na​2S​O3


5. In which compound is the percent by mass of oxygen the greatest?

a. KClO b. KClO​2​ c. KClO3​ ​ d. KClO​4


6. A solution is prepared by adding 20.0 g of NaNO​3​ to a beaker and then diluting to a total of 100 grams with water. What is the percent mass of NaNO​3​ in the solution?


7. A new company claims that its high energy drink contains 10% by mass of potassium ion. An independent laboratory buys a bottle containing 450 grams of solution. If the company is telling the truth, how much potassium should be in the solution?


8. The percent composition by mass of magnesium in MgS (gram-formula mass = 56 grams/mole) is equal to

(1) 56/32. Times 100

(2) 32./56 times 100

(3) 56/24 time 100

(4) 24/56 times 100


9. What is the percent composition by mass of hydrogen in NH​4H​ CO​3​ (gram-formula mass = 79 grams/mole)?

(1) 5.1% (3) 10.% (2) 6.3% ​ (4) 50.%



10. Calculate the percent composition of aluminum in CeMgAl​11O​19




1
Expert's answer
2021-01-28T05:36:10-0500

1.)carbon monoxide formula is CO

Composition of CO= "12.01+16.00=28.01"

"\%composition=\dfrac{12.01}{28.01}*100=57\%"

2.) Composition of"NaHCO_3"

\%composition of oxygen

="\dfrac{48}{84}*100=57.14\%"


3.) percentage of carbon in sucrose of "C_{12}H_{22}O_{11}"

"=\dfrac{144.12}{342.34}*100=42\%"

4.)

Composition of sodium in "Na_2SO_4=\dfrac{45.96}{142.02}*100=32.36\%"

Sodium composition in "Na_2SO_3=\dfrac{45.96}{128.06}*100= 35\%"

Na2SO3has the greatest composition of sodium.

5.)KClO3 has the greatest composition of Oxygen "=66.17\%"

7.)

"\dfrac{x}{450}*100=10\%\newline 100x=4500\newline potassium =45g"

8.) composition of magnesium in 56g/mole.

The correct answer is 4

"\dfrac{24.3}{56}*100"

9. Percentage composition of Hydrogen in "NH_4HCO_3"

"1.01*5=5.05\newline =\dfrac{5.05}{79}*100=6.67\%"

10.)

Percentage composition of aluminum